OpenStudy (nicole143):

Write an explicit formula for the geometric sequence. sequence: 25, 5, 1, 1/5 @mathmale Can you explain this?

OpenStudy (nicole143):

@kirbykirby Can you help?

OpenStudy (kirbykirby):

\[a_n=\frac{25}{5^n}\]

OpenStudy (nicole143):

Can you further explain?

OpenStudy (mathmale):

Nicole, please identify the first term of this sequence. Call it "a sub 1."\[a _{1}=?\]

OpenStudy (kirbykirby):

\[a_n=\frac{25}{5^n}\] for n=0,1,2,3

OpenStudy (nicole143):

Okay, so each next term is a division of 5?

OpenStudy (kirbykirby):

yes

OpenStudy (mathmale):

Now that you have the first term, how do you get the second? Hint: You multiply the first term by what? Each succeeding term is a fractional multiple of the previous term.

OpenStudy (mathmale):

Nicole, please identify the first term of this sequence. Call it "a sub 1."

OpenStudy (nicole143):

So whenever you have a sequence you find the common ration and put it under the first term?

OpenStudy (nicole143):

First term = 25

OpenStudy (kirbykirby):

I suppose you can also use this notation\[a_n=\left\{\frac{25}{5^n}\right\}_{n=0} ^3\]

OpenStudy (mathmale):

Yes, if we're talking about geometric sequences. Yes, the first term is 25. From that first term, we get the next term, 5, by multiplying 25 by what common ratio?

OpenStudy (nicole143):

So whenever you have a sequence you find the common ration and put it under the first term? To find the explicit formula..

OpenStudy (nicole143):

5 @ MM

OpenStudy (anonymous):

It is a G.P. first term=25 common ratio r=5/25=1/5 \[tn=ar ^{n-1}=25\left( \frac{ 1 }{ 5 } \right)^{n-1}\]

OpenStudy (mathmale):

Nicole, at this point the key is to find the common ratio. 25 times 5 is 125, not 5, so 5 could not be the common ratio. Hint" the common ratio is a fraction.

OpenStudy (nicole143):

Wait, I thought what Kirby said was the answer and what I said was how you got it..

OpenStudy (mathmale):

Surjithayer (and kirbykirby) have done this nicely, showing that the initial term is 25 and that the common ratio is 1/5.

OpenStudy (nicole143):

Okay

OpenStudy (nicole143):

That ones done right?

OpenStudy (kirbykirby):

Yeah I think mathmale is giving you a more general idea on how to find these formulae. The sequence given though looked trivial to me so I just write the formula. But yeah you are just dividing by 5 every time. But if you have a more complicated sequence, then you can follow mathmale's advice (particularly if there's an "added constant")

OpenStudy (mathmale):

Let's try explaining that in different words: "The first term is 25. The next term is 5. By what number (ratio or fraction) must we multiply the first term to obtain the second? You could write 25x=5, then solve for x: x=5/25 = 1/5. Again, the common ratio is 1/5.

OpenStudy (nicole143):

Okay, I understand that part now

OpenStudy (mathmale):

Both kirbykirby and Surjithayer show a starting value for n. If we use kirbykirby's model, then \[a _{n}=25(1/5)^n,\]with n starting at 0.

OpenStudy (nicole143):

Wait, um.. Find x.. 0.2?

OpenStudy (mathmale):

Actually, n goes to infinity, since the problem does not ask you to "find the first 3 or 4 terms". Note that 1/5 = 0.2; you could use either (but not both) in writing your formula for the nth term of the sequence.

OpenStudy (mathmale):

Summarizing, if we start with the geom. sequence 25, 5, 1, 1/5, 1/25, the first term, a, is 25; the common ratio, r, is 1/5, and n is merely a counter. Please look at both kirbykirby's and Surjithayer's formulas for a-sub-n and then choose whichever is clearer to you.

OpenStudy (nicole143):

Okay, thank you! @mathmale @kirbykirby Could you help me with another?

OpenStudy (mathmale):

Please post it, and, as before, I'll help you all I can. I'm helping a number of other students simultaneously. Nice work, Nicole.

OpenStudy (nicole143):

Okay!